I need to graph 3x+4y=20
I got the equation this far y=-3/4+5
I can't figure out how to graph it now

Answer by mananth(16946)

y=(-3/4)x+5

find the Y & X intercept points
plot the points and draw the line graph
let x=0, y = 5 (0,5)
Let x=4, y= 2(4,2)
Let y =0, x=20/3 ( 20/3, 0)